Problem 2: Let X and Y be two independent random variables with E(X)=1, E(Y)=0, Var(X)=4, Var(Y)=2. Let W=2X +Y+1 and Z=3X+Y.
a) Find the expected value of W, the expected value of Z.
b) Find the variance of W.
c) Find the covariance of Z and W.
1
Expert's answer
2019-11-25T12:00:51-0500

Given that "X" and "Y"are two independent random variables

"E(X)=1, E(Y)=0, Var(X)=4, Var(Y)=2""W=2X+Y+1, Z=3X+Y"

a) Distributive property of expected value: for any rv "X,Y" and any constants,"a,b,c"


"E(aX+bY+c)=aE(X)+b(Y)+c"

"E(W)=E(2X+Y+1)=2E(X)+E(Y)+1=""=2(1)+0+1=3"

"E(Z)=E(3X+Y)=3E(X)+E(Y)=""=3(1)+0=3"

b) If "X" is any random variable and "c" is any constant, then


"Var(cX)=c^2Var(X)""Var(X+c)=Var(X)"

If "X"and "Y"are two independent random variables, then


"Var(X+Y)=Var(X)+Var(Y)"


"Var(W)=Var(2X+Y+1)=4Var(X)+Var(Y)=""=4(4)+2=18"


c) For any two random variables "X" and "Y,"


"Cov(X, Y)=E(XY)-\mu_X\mu_Y""Cov(X,Y)=Cov(Y,X)""Cov(X,X)=Var(X)"

If "X" and "Y" are independent, then "Cov(X,Y)=0." by observing that "E(XY)=E(X)E(Y)"


Distributive property of covariance: for any rv "X,Y,Z" and any constants, a, b, c,


"Cov(aX+bY+c, Z)=aCov(X, Z)+bCov(Y,Z)"

"Cov(W,Z)=Cov(Z,W)=""=Cov(2X+Y+1,3X+Y)=""=2Cov(X,3X+Y)+Cov(Y,3X+Y)=""=6Cov(X,X)+2Cov(X,Y)+3Cov(Y,X)+Cov(Y,Y)=""=6Var(X)+5Cov(X,Y)+Var(Y)"

"Cov(W,Z)=6(4)+5(0)+2=26"